# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ Abstract
I Introduction
Boson Sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [1, 2, 3]. Accordingly, Boson Sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of Boson Sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [4, 5, 6, 7]. The problem is that those noises would possibly rule out the classical intractability of Boson Sampling, and thus potentially hinder the experimental demonstration of quantum advantage with Boson Sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] have proposed efficient classical simulation algorithms of Boson Sampling under various noise models, such as photon loss, partial distinguishability, random Gaussian noise, etc. Their results indicate that as the noise rate of boson sampler increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
[…]
@@ I Introduction
I.1 Outlines: average-case hardness of shallow-depth Boson Sampling
@@
Most of the current theoretical foundations of the average-case hardness of Boson Sampling rely on global Haar random unitary circuits [1, 2, 3, 27, 28, 29, 30], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem for our case is that implementing a global Haar random unitary requires at least polynomially large circuit depth (e.g., see [31, 32]), and thus is not implementable in sub-polynomial circuit depths. This poses a challenge in proving the average-case hardness of Boson Sampling in the shallow-depth regime—specifically, the lack of the hiding property. Due to the absence of hiding in the shallow-depth regime, the random outcome instances of Boson Sampling cannot be hidden by random circuit instances. Accordingly, even if we establish average-case hardness over randomly chosen circuits for a fixed outcome, it does not directly imply classical simulation hardness via Stockmeyer’s reduction, which requires average-case hardness over randomly chosen outcomes [1].
@@
Another problem is that there already exist efficient classical algorithms that can approximately simulate shallow-depth Boson Sampling in certain circumstances, which directly rule out the classical simulation hardness in the shallow-depth regime for those cases. While exact simulation of Boson Sampling is classically hard even for constant-depth circuits [33], approximate simulation becomes easy for 1-dimensional local log-depth circuits [34, 35] and also for more general dimension local circuits under some constraints [36, 37]. According to their results, if we use circuits composed of only geometrically local gates, a polynomial circuit depth is generally required for a sufficiently large correlation to achieve approximate simulation hardness. Those results indicate that we cannot expect the hardness results in the most general case of shallow-depth circuits composed of local gates only.
@@
To deal with those problems, we take the following approach: First, we consider shallow linear optical circuit architectures that fully employ geometrically non-local gates. In fact, the implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [38, 39] and photonic architecture [6] implemented long-range interactions. Because of the absence of hiding, we derive the average-case hardness over both randomly chosen outcomes and randomly chosen circuits over the shallow circuit architectures, by establishing a worst-to-average-case reduction for both outcome and circuit instances. In short, this reduction process can be done in two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over a local random circuit distribution.
@@
For random circuit ensembles, we employ a local random circuit ensemble in the shallow circuit architecture, inspired by the hardness results of random circuit sampling [40, 41, 28, 42, 43]. Here, the local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of Boson Sampling [4, 5, 6, 7] follow a similar circuit distribution, but with geometrically local architectures. Additionally, due to the absence of symmetry, the local random circuit ensemble itself does not guarantee the bosonic birthday paradox, which is crucial for the hardness of Boson Sampling when using only collision-free outcomes [1]. Hence, we also include a random permutation circuit at the input of our random circuit to guarantee the bosonic birthday paradox for our random circuit ensemble (see Appendix B). In fact, this random permutation circuit can be implemented within the shallow depth regime (see, e.g., Lemma 1), and can be experimentally realized by preparing a random input configuration instead.
@@
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of geometrically non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of Boson Sampling in the logarithmic depth regime, for photon number $N$ and mode number $M\propto N^{\gamma}$ with $\gamma\geq 2$.
@@
There exists an $O(\log N)$-depth linear optical circuit architecture $\mathcal{A}$ composed of $O(N^{\gamma}\log N)$ number of geometrically non-local gates, such that approximating output probability of Boson Sampling within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$ with high probability over the outcomes and circuits in $\mathcal{A}$ is #$\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
Also, since our average-case hardness result regards both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [1]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth Boson Sampling.
@@
For $M\propto N^{\gamma}$ with $\gamma\geq 2$, if the allowed additive imprecision for the problem in Theorem 1 to be #$\rm{P}$-hard is improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$, the approximate Boson Sampling for the shallow-depth circuit in Theorem 1, up to constant total variation distance, is classically hard to simulate.
Now we provide an outline of our results, which is depicted in Fig. 1. We first define in Sec. II a shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ composed of non-local gates, which we will use throughout our results. Next, in Sec. III, we prove the worst-case #P-hardness of approximating output probability $p_{\bm{s}}(C)$ of a fixed outcome $\bm{s}$ of Boson Sampling, for any circuit $C$ in the shallow circuit architecture previously defined. In Sec. IV, we prove the average-case #P-hardness of approximating output probability $p_{\bm{s}}(U)$ for randomly chosen outcome $\bm{s}$ and randomly chosen circuit $U$ in the shallow circuit architecture, by establishing worst-to-average-case reduction. We prove in Sec. V how our average-case hardness results over both the random outcomes and random circuits lead to the classical simulation hardness. We also extend our average-case hardness result to the Gaussian Boson Sampling scheme in Sec. VI, and to the lossy Boson Sampling subject to photon loss channels in Sec. VII. In Sec. VIII, we conclude with several remarks.
@@
II Our setting
II.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible $N$-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [1, 44] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing $N$-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an $N$-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
Hereafter, we use the terminology Boson Sampling representing Boson Sampling with Fock-state inputs as in the original Boson Sampling proposal [1]; we will explicitly use the terminology Gaussian Boson Sampling for Boson Sampling with Gaussian state inputs [2]. We now define $p_{\bm{s}}(C)$ as the output probability of Boson Sampling with a linear optical circuit (unitary) matrix $C$ to obtain an outcome $\bm{s}$ from a fixed input configuration $\bm{t}$. More precisely, $p_{\bm{s}}(C)$ can be represented as [1]
@@ II.1 Notations
where $C_{\bm{s},\bm{t}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$. Here, we omit the $\bm{t}$ dependence in the output probability $p_{\bm{s}}(C)$ for simplicity because we will fix the input configuration throughout the paper.
We note that an $M$-mode linear optical circuit can be represented by an $M$ by $M$ unitary matrix in U$(M)$, which unitarily transforms mode operators. Specifically, we can represent a single-mode gate (i.e., a phase shifter) as a U$(1)$ matrix to the mode, and a two-mode gate (i.e., a beam splitter) as a U$(2)$ matrix along the modes. Also, the parallel application of gates can be represented as a unitary matrix with a block matrix form, and the serial application of gates can be represented as matrix multiplication of the unitary matrices. Accordingly to this correspondence, throughout this work, we will interchangeably use the terminology ‘(linear optical) circuit’ and ‘(unitary) matrix’.
@@
In this work, we consider the shallow-depth circuit architecture, specifically in logarithmic depth regime $D=O(\log N)$. As we only consider the binary logarithm throughout this work, we will omit the base of the logarithm such that $\log(\cdot)$ implies $\log_{2}(\cdot)$ hereafter. We now define the shallow linear optical circuit architecture of circuit depth $D=\log M$ with the number of gates $m=\frac{M}{2}\log M$, using the convention used in [45, 46].
@@
We emphasize that in Definition 2, gates are applied in parallel for all $j=1,2,\dots,2^{D-L}$ and $k=1,2,\dots,2^{L-1}$. Hence, by definition of the circuit depth, the total depth for $\mathcal{B}^{*}$ (and thus also $\mathcal{B}^{*}$) is given by $D=\log M$. We illustrate in Fig. 2 the circuit architecture $\mathcal{B}$ and $\mathcal{B}^{*}$ for $162^{4}=$ 16 modes (such that the circuit depth is $D=4$ each), and the form of their corresponding unitary matrix.
Next, we define the Kaleidoscope circuit architecture proposed in [45], using the butterfly circuit architecture defined above.
@@
Here, the circuit depth of the $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth $D=O(\log N)$ for $q=O(1)$. Also, the number of gates for the $q$-Kaleidoscope architecture is $m=qM\log M=O(N^{\gamma}\log N)$ for $q=O(1)$. Note that, because all the gates are geometrically non-local except for the gates in the first and the last depth for each $\mathcal{B}\mathcal{B}^{*}$, the number of non-local gates also scales $O(N^{\gamma}\log N)$ for $q=O(1)$. Throughout this work, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. The motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis, which is that for $M$ a power of $2$, any $M$-mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$:
Lemma 1 (Dao et al [45]).
Let $\bm{P}$ be an arbitrary $M\times M$ permutation matrix with $M$ a power of 2. Then $\bm{P}$ can be efficiently implemented in $\mathcal{B}\mathcal{B}^{*}$ using $M\log M$ number of two-mode permutation gates, i.e., $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
II.2 Definitions of random circuit ensemble and random outcome ensemble

### v3

@@ Abstract
1 Introduction
1.1 Backgrounds and motivation
Boson Sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [aaronson2011computational, hamilton2017gaussian, deshpande2022quantum]. Accordingly, Boson Sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of Boson Sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian]. The problem is that those noises would possibly rule out the classical intractability of Boson Sampling, and thus potentially hinder the experimental demonstration of quantum advantage with Boson Sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [renema2018classical, renema2018efficient, moylett2019classically, renema2020simulability, shi2022effect, van2024efficient, shchesnovich2019noise, oszmaniec2018classical, garcia2019simulating, qi2020regimes, brod2020classical, villalonga2021efficient, oh2021classical2, bulmer2022boundary, liu2023simulating, oh2023classical, oh2024classicalalgorithm, oh2025classical, oh2025recent] have proposed efficient classical simulation algorithms of Boson Sampling under various noise models, such as photon loss, partial distinguishability, random Gaussian noise, etc. Their results indicate that as the noise rate of Boson Sampling increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
A viable alternative to preclude the classical simulability due to the inevitable noise is to consider Boson Sampling with shallow-depth linear optical circuits, where the noise rate can be highly reduced. Specifically, among the shallow-depth regime, our primary focus is on investigating the simulation hardness for $\it{logarithmic}$ depth circuits; the intuition behind investigating logarithmic depth circuits lies in the potential to offer a “sweet-spot" regime for the hardness of Boson Sampling. Namely, this depth regime may avoid significant increases in noise rates to prevent classical simulability, while still being sufficiently large to generate quantum correlations and uphold classical intractability. Despite such intuitive understanding, the hardness argument of Boson Sampling in this shallow-depth regime, particularly from a complexity-theoretical perspective, has been less studied so far and thus remains widely open. Hence, our goal is to establish the complexity-theoretical foundations of the classical hardness of shallow-depth Boson Sampling, to suppress the classical simulability by noise in a rigorous manner and obtain a more noise-tolerant demonstration of quantum advantage with Boson Sampling.
@@ 1.1 Backgrounds and motivation
1.2 Our results: Average-case hardness of shallow-depth Boson Sampling
@@
Most of the current theoretical foundations of the average-case hardness of Boson Sampling rely on global Haar random unitary circuits [aaronson2011computational, hamilton2017gaussian, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2023complexity, bouland2024average], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem for our case is that implementing a global Haar random unitary requires at least polynomially large circuit depth (e.g., see [zyczkowski1994random, russell2017direct]), and thus is not implementable in sub-polynomial circuit depths. This poses a challenge in proving the average-case hardness of Boson Sampling in the shallow-depth regime—specifically, the lack of the hiding property. Due to the absence of hiding in the shallow-depth regime, the random outcome instances of Boson Sampling cannot be hidden by random circuit instances. Accordingly, even if we establish average-case hardness over randomly chosen circuits for a fixed outcome, it does not directly imply classical simulation hardness via Stockmeyer’s reduction, which requires average-case hardness over randomly chosen outcomes [aaronson2011computational].
@@
Another problem is that there already exist efficient classical algorithms that can approximately simulate shallow-depth Boson Sampling in certain circumstances, which directly rule out the classical simulation hardness in the shallow-depth regime for those cases. While exact simulation of Boson Sampling is classically hard even for constant-depth circuits [brod2015complexity], approximate simulation becomes easy for 1-dimensional local log-depth circuits [vidal2003efficient, qi2022efficient] and also for more general dimension local circuits under some constraints [deshpande2018dynamical, oh2022classical, kolarovszki2023simulating]. According to their results, if we use circuits composed of only geometrically local gates, a polynomial circuit depth is generally required for a sufficiently large correlation to achieve approximate simulation hardness. Those results indicate that we cannot expect the hardness results in the most general case of shallow-depth circuits composed of local gates only.
@@
To deal with those problems, we take the following approach: First, we consider shallow linear optical circuit architectures that fully employ geometrically non-local gates. In fact, the implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [shen2014scalable, chen2023scalable] and photonic architecture [madsen2022quantum] implemented long-range interactions. Because of the absence of hiding, we derive the average-case hardness over both randomly chosen outcomes and randomly chosen circuits over the shallow circuit architectures, by establishing a worst-to-average-case reduction for both outcome and circuit instances. In short, this reduction process can be done in two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over a local random circuit distribution.
@@
For random circuit ensembles, we employ a local random circuit ensemble in the shallow circuit architecture, inspired by the hardness results of random circuit sampling [bouland2019complexity, movassagh2023hardness, bouland2022noise, kondo2022quantum, krovi2022average]. Here, the local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of Boson Sampling [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] follow a similar circuit distribution, but with geometrically local architectures. Additionally, due to the absence of symmetry, the local random circuit ensemble itself does not guarantee the bosonic birthday paradox, which is crucial for the hardness of Boson Sampling when using only collision-free outcomes [aaronson2011computational]. Hence, we also include a random permutation circuit at the input of our random circuit to guarantee the bosonic birthday paradox for our random circuit ensemble (see Appendix B). In fact, this random permutation circuit can be implemented within the shallow depth regime (see, e.g., Lemma 1), and can be experimentally realized by preparing a random input configuration instead.
@@
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of geometrically non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of Boson Sampling in the logarithmic depth regime, for photon number NN and mode number $M\propto N^{\gamma}$ with $\gamma\geq 2$.
@@
Consider a Boson Sampling problem with NN photons over $M=\Omega(N^{\gamma})$ mode, where $\gamma\geq 2$. There exists an $O(\log N)$-depth linear optical circuit architecture $\mathcal{A}$, consisting of $O(N^{\gamma}\log N)$ number of geometrically non-local gates, such that estimating the output probability to within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$, with high probability over the choice of outcomes and circuits in $\mathcal{A}$, is $\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
Also, since our average-case hardness result regards both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [aaronson2011computational]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth Boson Sampling.
@@
For $M\propto N^{\gamma}$ with $\gamma\geq 2$, if the allowed additive imprecision for the problem in Theorem 1 to be $\rm{P}$-hard is improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$, the approximate Boson Sampling for the shallow-depth circuit in Theorem 1, up to constant total variation distance, is classically hard to simulate.
1.3 Paper organization
Our paper is organized as follows, which is outlined in Fig. 1. We first define in Sec. 2 a shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ composed of non-local gates, which we will use throughout our results. Next, in Sec. 3, we prove the worst-case #P-hardness of approximating output probability $p_{\bm{s}}(C)$ of a fixed outcome $\bm{s}$ of Boson Sampling, for any circuit $C$ in the shallow circuit architecture previously defined. In Sec. 4, we prove the average-case #P-hardness of approximating output probability $p_{\bm{s}}(U)$ for randomly chosen outcome $\bm{s}$ and randomly chosen circuit $U$ in the shallow circuit architecture, by establishing worst-to-average-case reduction. We prove in Sec. 5 how our average-case hardness results over both the random outcomes and random circuits lead to the classical simulation hardness. We also extend our average-case hardness result to the Gaussian Boson Sampling scheme in Sec. 6, and to the lossy Boson Sampling subject to photon loss channels in Sec. 7. In Sec. 8, we conclude with several remarks.
@@ 1.3 Paper organization
2 Our settings
2.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number NN polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible NN-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [aaronson2011computational, arkhipov2012bosonic] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing NN-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an NN-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
Hereafter, we use the terminology Boson Sampling representing Boson Sampling with Fock-state inputs as in the original Boson Sampling proposal [aaronson2011computational]; we will explicitly use the terminology Gaussian Boson Sampling for Boson Sampling with Gaussian state inputs [hamilton2017gaussian]. We now define $p_{\bm{s}}(C)$ as the output probability of Boson Sampling with a linear optical circuit (unitary) matrix $C$ to obtain an outcome $\bm{s}$ from a fixed input configuration $\bm{t}$. More precisely, $p_{\bm{s}}(C)$ can be represented as [aaronson2011computational]
@@ 2.1 Notations
where $C_{\bm{s},\bm{t}}$ is an NN by NN matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$. Here, we omit the $\bm{t}$ dependence in the output probability $p_{\bm{s}}(C)$ for simplicity because we will fix the input configuration throughout the paper.
We note that an $M$-mode linear optical circuit can be represented by an $M$ by $M$ unitary matrix in U(M)(M), which unitarily transforms mode operators. Specifically, we can represent a single-mode gate (i.e., a phase shifter) as a U(1)(1) matrix to the mode, and a two-mode gate (i.e., a beam splitter) as a U(2)(2) matrix along the modes. Also, the parallel application of gates can be represented as a unitary matrix with a block matrix form, and the serial application of gates can be represented as matrix multiplication of the unitary matrices. Accordingly to this correspondence, throughout this work, we will interchangeably use the terminology ‘(linear optical) circuit’ and ‘(unitary) matrix’.
@@
In this work, we consider the shallow-depth circuit architecture, specifically in logarithmic depth regime $D=O(\log N)$. As we only consider the binary logarithm throughout this work, we will omit the base of the logarithm such that $\log(\cdot)$ implies $\log_{2}(\cdot)$ hereafter. We now define the shallow linear optical circuit architecture of circuit depth $D=\log M$ with the number of gates $m=\frac{M}{2}\log M$, using the convention used in [dao2020kaleidoscope, go2024exploring].
@@
We emphasize that in Definition 2, gates are applied in parallel for all $j=1,2,\dots,2^{D-L}$ and $k=1,2,\dots,2^{L-1}$. Hence, by definition of the circuit depth, the total depth for $\mathcal{B}^{*}$ (and thus also $\mathcal{B}^{*}$) is given by $D=\log M$. We illustrate in Fig. 2 the circuit architecture $\mathcal{B}$ and $\mathcal{B}^{*}$ for $2^{4}=16$ modes (such that the circuit depth is $D=4$ each), and the form of their corresponding unitary matrix.
Next, we define the Kaleidoscope circuit architecture proposed in [dao2020kaleidoscope], using the butterfly circuit architecture defined above.
@@
Here, the circuit depth of the $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth $D=O(\log N)$ for $q=O(1)$. Also, the number of gates for the $q$-Kaleidoscope architecture is $m=qM\log M=O(N^{\gamma}\log N)$ for $q=O(1)$. Note that, because all the gates are geometrically non-local except for the gates in the first and the last depth for each $\mathcal{B}\mathcal{B}^{*}$, the number of non-local gates also scales $O(N^{\gamma}\log N)$ for $q=O(1)$. Throughout this work, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. The motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis, which is that for $M$ a power of 22, any $M$-mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$:
Lemma 1 (Dao et al [dao2020kaleidoscope]).
Let $\bm{P}$ be an arbitrary $M\times M$ permutation matrix with $M$ a power of 2. Then $\bm{P}$ can be efficiently implemented in $\mathcal{B}\mathcal{B}^{*}$ using $M\log M$-number of two-mode permutation gates, i.e., $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
2.2 Definitions for random circuit and outcome ensemble